# Asymptotic theory of in-context learning by linear attention

### v2

@@ Acknowledgments
SI.6.1 Ridge estimator error
@@
$y_{i}=x_{i}^{\top}w+\epsilon_{i},\quad\text{for }i=1,2,\ldots,\ell,$ (SI.6.1)
@@
$\mathbb{E}[ww^{\top}]=I_{d},\qquad\mathbb{E}[yw^{\top}]=X^{\top},\qquad\text{and }\mathbb{E}[yy^{\top}]=X^{\top}X+\rho I_{\ell},$ (SI.6.2)
@@
$\beta=(XX^{\top}+\rho I_{d})^{-1}Xy+Z,$ (SI.6.3)
where $Z$ is a centered Gaussian vector independent of $y$. Note that the matrix in front of $y$ in the above expression is chosen so that the cross-correlation $\mathbb{E}[yw^{\top}]$ matches that in (SI.6.2). It then immediately follows that the Bayesian estimator of $w$ is
$w^{\text{ridge}}_{\text{Bayes}}=(XX^{\top}+\rho I_{d})^{-1}Xy.$ (SI.6.4)
@@
$\hat{y}_{\ell+1}$ $=x_{\ell+1}^{\top}w^{\text{ridge}}_{\text{Bayes}},$ (SI.6.5)
@@
$e_{\text{Bayesian}}$ $=\mathbb{E}[(y_{\ell+1}-x_{\ell+1}^{\top}w^{\text{ridge}}_{\text{Bayes}})]$ (SI.6.6)
$=\rho+\frac{1}{d}\mathbb{E}\left[\lVert w-w^{\text{ridge}}_{\text{Bayes}}\rVert^{2}\right]$ (SI.6.7)
$\simeq\rho\left(1+\frac{1}{\alpha}\mathcal{M}_{\alpha}\left(\frac{\rho}{\alpha(1-\gamma)}\right)\right),$ (SI.6.8)
where $\mathcal{M}_{\alpha}(x)$ is the function defined in (SI.8.4). In the large $\alpha$ limit, we have
$e_{\text{Bayesian}}\simeq\rho(1+1/\alpha).$ (SI.6.9)
SI.6.2 Asymptotics of the linear transformer in $\kappa$
We wish to expand (SI.5.40) and (SI.5.51) in large $\kappa$ to study the decay of $g_{\text{task}}=e_{\text{ICL}}-e_{\text{IDG}}$ toward 0 as $\kappa\to\infty$. By expanding auxillary variables in $\kappa$, we have
$\mathcal{M}_{\kappa}(q)\approx\frac{1}{1+q}\left(1+\frac{1}{(1+q)^{2}\kappa}\right)\,,\qquad\mathcal{M}^{\prime}_{\kappa}(q)\approx-\frac{1}{(1+q)^{2}}\left(1+\frac{3}{(1+q)^{2}\kappa}\right)$
@@
$\mu=q\mathcal{M}_{\kappa/\tau}(q)\approx\frac{q}{1+q}\left(1+\frac{\tau}{(1+q)^{2}\kappa}\right)\,,\qquad\xi\approx\frac{1-\tau}{\tau}(1+q)\left(1-\frac{\tau}{(1+q)^{2}\kappa}\right)\,,\qquad p\approx 1+\frac{\tau^{2}}{(1+\tau)^{2}\kappa}\,.$
Substituting these simplifications into (SI.5.40) and (SI.5.51) gives the desired characterization
@@
$c_{1}$ $\equiv\frac{2\left(\left(-2q-\frac{1}{2}\right)\tau^{2}+\left(q^{2}+\left(-\rho+3\right)q+1-\rho\right)\tau-\left(1+q\right)\left(q+\frac{1}{2}-\frac{\rho}{2}\right)\right)\tau}{\left(-1+\tau\right)\left(1+q\right)^{3}}\,,$ (SI.6.10)
$c_{2}$ $\equiv 2\frac{q^{2}}{(1+q)^{3}}+\frac{1}{(\tau-1)(q+1)^{2}}\left(\rho+q-\frac{q^{2}}{q+1}\right)\,.$ (SI.6.11)
SI.6.3 Asymptotics of the dMMSE estimator in $\kappa$
@@
$g^{\text{dMMSE}}_{\text{task}}$ $=\frac{1}{d}\mathbb{E}_{w^{\ast}\sim\mathcal{P}_{\text{test}}}\left[\mathbb{E}_{w_{i}\sim\mathcal{P}_{\text{train}}}\left[\min_{i\in[k]}\|w^{\ast}-w_{i}\|^{2}\right]\right].$ (SI.6.12)
@@
$\|w^{\ast}-w_{i}\|^{2}$ $=2d-2\sqrt{d}x_{i}\qquad\text{ for $x_{i}$ a single component of vector $w_{i}\sim\text{Unif}(\mathcal{S}^{d-1}(\sqrt{d}))$}$
$=2d-2\sqrt{d}(2\sqrt{d}\,b_{i}-\sqrt{d})\qquad\text{for $b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)$}$
$=4d-4db_{i}$
@@
$g_{\text{task}}$ $=\rho+4-4\,\mathbb{E}_{b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)}\left[\max_{i\in[k]}b_{i}\right].$
@@
$\mathbb{E}_{b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)}\left[\max_{i\in[k]}b_{i}\right],$ (SI.6.13)
@@
$B(D)=\frac{\Gamma(D)^{2}}{\Gamma(2D)}.$ (SI.6.15)
@@
[…]
@@
[…]
Naively, it looks like one could get a simple series expansion in $1/k$ through integration by parts, but this will fail because $\left.\frac{dM}{dt}\right|_{t=0}=\infty$ for $D>1$. One can see this by differentiating (SI.6.20) with respect to $t$. Instead, have
[…]
@@
[…]
To get the leading behavior, one proceeds by expanding in leading order to $1-M$ or the right and to $t/k$ on the left
$1-\frac{t}{k}=1-\frac{\epsilon(t/k)^{D}}{B(D)}\quad\implies\epsilon\approx\left(\frac{B(D)t}{k}\right)^{1/D}$ (SI.6.25)
@@
SI.7 Equivalent Statistical Representations
In this appendix, we present an equivalent (but simplified) statistical model for the regression vector $H_{Z}$. This statistica lly-equivalent model will simplify the moment calculations in Section SI.2 and the random matrix analysis in Section SI.4.
@@
Let $w$ be a given task vector with $\lVert w\rVert=\sqrt{d}$. Meanwhile, let $a\sim\mathcal{N}(0,1)$, $s\sim\mathcal{N}(0,1)$, $\epsilon\sim\mathcal{N}(0,\rho)$ be three scalar normal random variables, and $q\sim\mathcal{N}(0,I_{\ell-1})$, $g\sim\mathcal{N}(0,I_{d-1})$, $u\sim\mathcal{N}(0,I_{d-1})$, and $v_{\epsilon}\sim\mathcal{N}(0,\rho I_{\ell})$ be isotropic normal random vectors. Moreover, $w$ and all of the above random variables are mutually independent. We have the following equivalent statistical representation of the pair $(H_{Z},y_{\ell+1})$:
$H_{Z}\overset{(d)}{=}(d/\ell)M_{w}\begin{bmatrix}s\\ u\end{bmatrix}\begin{bmatrix}h^{\top}M_{w},&(a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\end{bmatrix},$ (SI.7.1)
@@
$(M_{w})e_{1}=\frac{w}{\lVert w\rVert},$ (SI.7.3)
@@
$h\coloneqq\begin{bmatrix}\frac{{\theta_{\epsilon}}a}{\sqrt{d}}+\frac{a^{2}}{d}+\theta_{q}^{2}\\ \big{[}(\theta_{\epsilon}+a/\sqrt{d})^{2}+\theta_{q}^{2}\big{]}^{1/2}{g}/{\sqrt{d}}\end{bmatrix};$ (SI.7.4)
@@
$\theta_{\epsilon}={\lVert v_{\epsilon}\rVert}/{\sqrt{d}}\qquad\text{and}\qquad\theta_{q}=\lVert q\rVert/\sqrt{d}.$ (SI.7.5)
@@
For two random variables $A$ and $B$, the notation $A\overset{(d)}{=}B$ indicates that $A$ and $B$ have identical probability distributions. Note that $A$ and $B$ can be either scalars [as in the case of (SI.7.2)], or matrices of matching dimensions [as in the case of (SI.7.1)].
@@
A concrete construction of the symmetric and orthonormal matrix $M_{w}$ satisfying (SI.7.3) can be based on the Householder transformation [60, 61, 62].
@@
$x_{\ell+1}\overset{(d)}{=}\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}s\\ u\end{bmatrix},$ (SI.7.6)
where $s\sim\mathcal{N}(0,1)$ and $u\sim\mathcal{N}(0,I_{d-1})$ are two independent normal random variables (vectors), and $M_{w}$ is the symmetric orthonormal matrix specified in (SI.7.3). Note that $y_{\ell+1}=x_{\ell+1}^{\top}w+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\rho)$ denoting the noise. The representation in (SI.7.2) then follows immediately from (SI.7.6) and the identity in (SI.7.3).
To show (SI.7.1), we first reparameterize the $d\times\ell$ Gaussian data matrix $X$ as
$X=M_{w}\begin{bmatrix}a&q^{\top}\\ p&U\end{bmatrix}M_{v_{\epsilon}}/\sqrt{d}.$ (SI.7.7)
@@
$M_{v_{\epsilon}}e_{1}=\frac{v_{\epsilon}}{\lVert v_{\epsilon}\rVert},$ (SI.7.8)
where $e_{1}$ denotes the first natural basis vector in $\mathbb{R}^{\ell}$. Since the data matrix $X$, the task vector $w$, and the noise vector $v_{\epsilon}$ are mutually independent, it is straightforward to verify via the rotational symmetry of the isotropic normal distribution that both sides of (SI.7.7) have identical probability distributions. Using this new representation, we have
$Xv_{\epsilon}=\theta_{\epsilon}M_{w}\begin{bmatrix}a\\ p\end{bmatrix}.$ (SI.7.9)
@@
$X^{\top}w=M_{v_{\epsilon}}\begin{bmatrix}a\\ q\end{bmatrix},$ (SI.7.10)
@@
$XX^{\top}w=\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}a^{2}+\lVert q\rVert^{2}\\ ap+Uq\end{bmatrix}.$ (SI.7.11)
Combining (SI.7.10) and (SI.7.11) yields
$Xy$ $=XX^{\top}w+Xv_{\epsilon}$ (SI.7.12)
$=M_{w}\begin{bmatrix}\theta_{\epsilon}a+a^{2}/\sqrt{d}+\theta_{q}^{2}\sqrt{d}\\ (\theta_{\epsilon}+a/\sqrt{d})p+Uq/\sqrt{d}\end{bmatrix}.$ (SI.7.13)
@@
$\frac{1}{\sqrt{d}}Xy\overset{(d)}{=}M_{w}h,$ (SI.7.14)
where $h$ is the random vector defined in (SI.7.4).
@@
$y^{\top}y$ $=\lVert X^{\top}w+v_{\epsilon}\rVert^{2}$ (SI.7.15)
$=\lVert X^{\top}w+\theta_{\epsilon}\sqrt{d}M_{v_{\epsilon}}e_{1}\rVert^{2}$ (SI.7.16)
$=(a+\theta_{\epsilon}\sqrt{d})^{2}+\theta_{q}^{2}d,$ (SI.7.17)
where the second equality follows from (SI.7.8) and to reach the last equality we have used the representation in (SI.7.10). To show (SI.7.1), we recall the definition of $H_{Z}$ in (17). Substituting (SI.7.6), (SI.7.14) and (SI.7.17) into (17), we are done. ∎
SI.8 The Stieltjes Transforms of Wishart Ensembles

### v3

@@ Appendix F Task Generalization and Bayesian Estimators
F.1 Ridge estimator error
@@
$y_{i}=x_{i}^{\top}w+\epsilon_{i},\quad\text{for }i=1,2,\ldots,\ell,$ (201)
@@
$\mathbb{E}[ww^{\top}]=I_{d},\qquad\mathbb{E}[yw^{\top}]=X^{\top},\qquad\text{and }\mathbb{E}[yy^{\top}]=X^{\top}X+\rho I_{\ell},$ (202)
@@
$\beta=(XX^{\top}+\rho I_{d})^{-1}Xy+Z,$ (203)
where $Z$ is a centered Gaussian vector independent of $y$. Note that the matrix in front of $y$ in the above expression is chosen so that the cross-correlation $\mathbb{E}[yw^{\top}]$ matches that in (202). It then immediately follows that the Bayesian estimator of $w$ is
$w^{\text{ridge}}_{\text{Bayes}}=(XX^{\top}+\rho I_{d})^{-1}Xy.$ (204)
@@
$\hat{y}_{\ell+1}$ $=x_{\ell+1}^{\top}w^{\text{ridge}}_{\text{Bayes}},$ (205)
@@
$e_{\text{Bayesian}}$ $=\mathbb{E}[(y_{\ell+1}-x_{\ell+1}^{\top}w^{\text{ridge}}_{\text{Bayes}})]$ (206)
$=\rho+\frac{1}{d}\mathbb{E}\left[\mathinner{\!\left\lVert w-w^{\text{ridge}}_{\text{Bayes}}\right\rVert}^{2}\right]$ (207)
$\simeq\rho\left(1+\frac{1}{\alpha}\mathcal{M}_{\alpha}\left(\frac{\rho}{\alpha(1-\gamma)}\right)\right),$ (208)
where $\mathcal{M}_{\alpha}(x)$ is the function defined in (253). In the large $\alpha$ limit, we have
$e_{\text{Bayesian}}\simeq\rho(1+1/\alpha).$ (209)
F.2 Asymptotics of the linear transformer in $\kappa$
We wish to expand (186) and (197) in large $\kappa$ to study the decay of $g_{\text{task}}=e_{\text{ICL}}-e_{\text{IDG}}$ toward 0 as $\kappa\to\infty$. By expanding auxillary variables in $\kappa$, we have
$\mathcal{M}_{\kappa}(q)\approx\frac{1}{1+q}\left(1+\frac{1}{(1+q)^{2}\kappa}\right)\,,\qquad\mathcal{M}^{\prime}_{\kappa}(q)\approx-\frac{1}{(1+q)^{2}}\left(1+\frac{3}{(1+q)^{2}\kappa}\right)$ (210)
@@ F.2 Asymptotics of the linear transformer in $\kappa$
$\mu=q\mathcal{M}_{\kappa/\tau}(q)\approx\frac{q}{1+q}\left(1+\frac{\tau}{(1+q)^{2}\kappa}\right)\,,\qquad\xi\approx\frac{1-\tau}{\tau}(1+q)\left(1-\frac{\tau}{(1+q)^{2}\kappa}\right)\,,\qquad p\approx 1+\frac{\tau^{2}}{(1+\tau)^{2}\kappa}\,.$ (211)
Substituting these simplifications into (186) and (197) gives the desired characterization
@@
$c_{1}$ $\equiv\frac{2\left(\left(-2q-\frac{1}{2}\right)\tau^{2}+\left(q^{2}+\left(-\rho+3\right)q+1-\rho\right)\tau-\left(1+q\right)\left(q+\frac{1}{2}-\frac{\rho}{2}\right)\right)\tau}{\left(-1+\tau\right)\left(1+q\right)^{3}}\,$, (212)
$c_{2}$ $\equiv 2\frac{q^{2}}{(1+q)^{3}}+\frac{1}{(\tau-1)(q+1)^{2}}\left(\rho+q-\frac{q^{2}}{q+1}\right)\,.$ (213)
F.3 Asymptotics of the dMMSE estimator in $\kappa$
@@
$g^{\text{dMMSE}}_{\text{task}}$ $=\frac{1}{d}\mathbb{E}_{w^{\ast}\sim\mathcal{P}_{\text{test}}}\left[\mathbb{E}_{w_{i}\sim\mathcal{P}_{\text{train}}}\left[\min_{i\in[k]}\|w^{\ast}-w_{i}\|^{2}\right]\right].$ (214)
@@
[…]
$=2d-2\sqrt{d}(2\sqrt{d}\,b_{i}-\sqrt{d})\qquad\text{for $b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)$}$ (216)
$=4d-4db_{i}$ (217)
@@
$g_{\text{task}}$ $=\rho+4-4\,\mathbb{E}_{b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)}\left[\max_{i\in[k]}b_{i}\right].$ (218)
@@
$\mathbb{E}_{b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)}\left[\max_{i\in[k]}b_{i}\right],$ (219)
@@
$B(D)=\frac{\Gamma(D)^{2}}{\Gamma(2D)}.$ (221)
@@
$\mathbb{E}_{b_{i}\sim\text{Beta}\left(\frac{d-1}{2},\frac{d-1}{2}\right)}\left[\max_{i\in[k]}b_{i}\right]=\frac{k}{B(D)}\int_{0}^{1}dMM^{D}(1-M)^{D-1}\left(\frac{1}{B(D)}\int_{0}^{M}dx\,(1-x)^{D-1}x^{D-1}\right)^{k-1}$ (224)
@@
$k\int_{0}^{\infty}dt\,M(t)e^{-kt}.$ (227)
Naively, it looks like one could get a simple series expansion in $1/k$ through integration by parts, but this will fail because $\left.\frac{dM}{dt}\right|_{t=0}=\infty$ for $D>1$. One can see this by differentiating (226) with respect to $t$. Instead, have
$k\int_{0}^{\infty}dt\,M(t)e^{-kt}=\int_{0}^{\infty}dt\,M(t/k)e^{-t}=1-\int_{0}^{\infty}dt\,\epsilon(t/k)e^{-t}$ (228)
@@
$e^{-t/k}$ $=\frac{1}{B(D)}\int_{0}^{1-\epsilon(t/k)}dx\,(1-x)^{D-1}x^{D-1}$ (230)
To get the leading behavior, one proceeds by expanding in leading order to 1M1-M or the right and to $t/k$ on the left
$1-\frac{t}{k}=1-\frac{\epsilon(t/k)^{D}}{B(D)}\quad\implies\epsilon\approx\left(\frac{B(D)t}{k}\right)^{1/D}$ (231)
@@
Appendix G Equivalent Statistical Representations
In this appendix, we present an equivalent (but simplified) statistical model for the regression vector $H_{Z}$. This statistically-equivalent model will simplify the moment calculations in Section B and the random matrix analysis in Section D.
@@ Appendix G Equivalent Statistical Representations
Let $w$ be a given task vector with $\mathinner{\!\left\lVert w\right\rVert}=\sqrt{d}$. Meanwhile, let $a\sim\mathcal{N}(0,1)$, $s\sim\mathcal{N}(0,1)$, $\epsilon\sim\mathcal{N}(0,\rho)$ be three scalar normal random variables, and $q\sim\mathcal{N}(0,I_{\ell-1})$, $g\sim\mathcal{N}(0,I_{d-1})$, $u\sim\mathcal{N}(0,I_{d-1})$, and $v_{\epsilon}\sim\mathcal{N}(0,\rho I_{\ell})$ be isotropic normal random vectors. Moreover, $w$ and all of the above random variables are mutually independent. We have the following equivalent statistical representation of the pair $(H_{Z},y_{\ell+1})$:
$H_{Z}\overset{(d)}{=}(d/\ell)M_{w}\begin{bmatrix}s\\ u\end{bmatrix}\begin{bmatrix}h^{\top}M_{w},&(a/\sqrt{d}+\theta_{\epsilon})^{2}/\sqrt{d}+\theta_{q}^{2}/\sqrt{d}\end{bmatrix}$, (233)
@@
$(M_{w})e_{1}=\frac{w}{\mathinner{\!\left\lVert w\right\rVert}},$ (235)
@@
$h\coloneqq\begin{bmatrix}\frac{{\theta_{\epsilon}}a}{\sqrt{d}}+\frac{a^{2}}{d}+\theta_{q}^{2}\\ \big[(\theta_{\epsilon}+a/\sqrt{d})^{2}+\theta_{q}^{2}\big]^{1/2}{g}/{\sqrt{d}}\end{bmatrix}$; (236)
@@
$\theta_{\epsilon}={\mathinner{\!\left\lVert v_{\epsilon}\right\rVert}}/{\sqrt{d}}\qquad\text{and}\qquad\theta_{q}=\mathinner{\!\left\lVert q\right\rVert}/\sqrt{d}.$ (237)
@@
For two random variables AA and $B$, the notation $A\overset{(d)}{=}B$ indicates that $A$ and $B$ have identical probability distributions. Note that $A$ and $B$ can be either scalars [as in the case of (234)], or matrices of matching dimensions [as in the case of (233)].
@@
A concrete construction of the symmetric and orthonormal matrix $M_{w}$ satisfying (235) can be based on the Householder transformation [60, 61, 62].
@@
$x_{\ell+1}\overset{(d)}{=}\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}s\\ u\end{bmatrix}$, (238)
where $s\sim\mathcal{N}(0,1)$ and $u\sim\mathcal{N}(0,I_{d-1})$ are two independent normal random variables (vectors), and $M_{w}$ is the symmetric orthonormal matrix specified in (235). Note that $y_{\ell+1}=x_{\ell+1}^{\top}w+\epsilon$, with $\epsilon\sim\mathcal{N}(0,\rho)$ denoting the noise. The representation in (234) then follows immediately from (238) and the identity in (235).
To show (233), we first reparameterize the $d\times\ell$ Gaussian data matrix $X$ as
$X=M_{w}\begin{bmatrix}a&q^{\top}\\ p&U\end{bmatrix}M_{v_{\epsilon}}/\sqrt{d}.$ (239)
@@
$M_{v_{\epsilon}}e_{1}=\frac{v_{\epsilon}}{\mathinner{\!\left\lVert v_{\epsilon}\right\rVert}},$ (240)
where $e_{1}$ denotes the first natural basis vector in $\mathbb{R}^{\ell}$. Since the data matrix $X$, the task vector $w$, and the noise vector $v_{\epsilon}$ are mutually independent, it is straightforward to verify via the rotational symmetry of the isotropic normal distribution that both sides of (239) have identical probability distributions. Using this new representation, we have
$Xv_{\epsilon}=\theta_{\epsilon}M_{w}\begin{bmatrix}a\\ p\end{bmatrix}$. (241)
@@
$X^{\top}w=M_{v_{\epsilon}}\begin{bmatrix}a\\ q\end{bmatrix}$, (242)
@@
$XX^{\top}w=\frac{1}{\sqrt{d}}M_{w}\begin{bmatrix}a^{2}+\mathinner{\!\left\lVert q\right\rVert}^{2}\\ ap+Uq\end{bmatrix}$. (243)
Combining (242) and (243) yields
$Xy$ $=XX^{\top}w+Xv_{\epsilon}$ (244)
$=M_{w}\begin{bmatrix}\theta_{\epsilon}a+a^{2}/\sqrt{d}+\theta_{q}^{2}\sqrt{d}\\ (\theta_{\epsilon}+a/\sqrt{d})p+Uq/\sqrt{d}\end{bmatrix}$. (245)
@@
$\frac{1}{\sqrt{d}}Xy\overset{(d)}{=}M_{w}h,$ (246)
where $h$ is the random vector defined in (236).
@@
$y^{\top}y$ $=\mathinner{\!\left\lVert X^{\top}w+v_{\epsilon}\right\rVert}^{2}$ (247)
$=\mathinner{\!\left\lVert X^{\top}w+\theta_{\epsilon}\sqrt{d}M_{v_{\epsilon}}e_{1}\right\rVert}^{2}$ (248)
$=(a+\theta_{\epsilon}\sqrt{d})^{2}+\theta_{q}^{2}d,$ (249)
where the second equality follows from (240) and to reach the last equality we have used the representation in (242). To show (233), we recall the definition of $H_{Z}$ in (65). Substituting (238), (246) and (249) into (65), we are done. ∎
Appendix H The Stieltjes Transforms of Wishart Ensembles